OpenStudy (albert0898):

SAT Math Level 2 Question The answer is D, but how do you get it? I know that 1 - cos = sin

OpenStudy (boldjon):

yeah what is it :3

OpenStudy (albert0898):

If \[\frac{ 1-\cos \theta }{ \sin \theta } = \frac{ \sqrt{3} }{ 3 }\] then, theta = (A) 15 (B) 30 (C) 45 (D) 60 (E) 75

OpenStudy (123ab456c):

C

OpenStudy (albert0898):

@123AB456C Wrong.

OpenStudy (hossamhoussien):

This link provides solve with full steps for your question look at it :) http://www.wolframalpha.com/input/?i=%7B%281-cos%28x%29%29%2Fsin%28x%29+%3D+sqrt%283%29%2F3%2C+0%3C%3D+x%2C+x+%3C%3D2pi%7D

OpenStudy (hossamhoussien):

the answer is 60 which is choice C

OpenStudy (baru):

like your multiplication tables, u need to memorize the values for cos and sin for the angles 0,30,60,90

OpenStudy (baru):

add 45 degrees also to that list :)

OpenStudy (hossamhoussien):

@Albert0898 i guess you didn't look at the link i posted earlier , right?

OpenStudy (albert0898):

@HossamHoussien I did, but I still don't get how that equation simplified. Although it's the right answer, I have no idea how the steps were computed. I've went through every rule and can't seem to simplify it the way Wolfram did.

OpenStudy (hossamhoussien):

Firstly you should notice that \[\frac{ \sqrt{3} }{ 3 }=\frac{ 1 }{ \sqrt{3} }\] and according to "Half-angel formula" it says that : \[\tan (\frac{ x }{ 2 })=\frac{ 1-\cos(\theta) }{ \sin(\theta) }\] So: \[\tan (\frac{ x }{ 2 }) = \frac{ \sqrt{3} }{ 3 }\] But you have to find value of (x) not (x/2) By calculator \[\frac{ x }{ 2 }=\tan^{-1} (\frac{ 1 }{ \sqrt3 })\] \[\frac{ x }{ 2 }=30\] \[x=60\] I hope that help